OpenStudy (anonymous):

Consider the following. f(x)=18x-15/-x^2+1 Determine the value that the function f approaches as the magnitude of x increases. (If you need to use or –, enter INFINITY or –INFINITY, respectively.)

OpenStudy (anonymous):

anybody can help me?

OpenStudy (amistre64):

the degree of the bottom is larger than the top; so it all goes to .... 0. i was wrong

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

is like 1/2,

OpenStudy (amistre64):

18x-15 ------- simple answer is divide everything by the highest x in the bottom -x^2+1

OpenStudy (anonymous):

you dont have to devide, you only divide when the numerador is bigger than the denominator. 2x^2/x for example

OpenStudy (amistre64):

\[\frac{\frac{18x}{x^2}-{15\over x^2}}{{-x^2\over x^2} + {1\over x^2}}\]

OpenStudy (anonymous):

but the answer is not infity.

OpenStudy (amistre64):

\[{(18/x) - (15/x^2)} = 0\]when x is large

OpenStudy (amistre64):

the answer is 0

OpenStudy (anonymous):

ut how you get 0?

OpenStudy (amistre64):

the concept is tp mulitply everything by a useful form of 1; I choose \(\frac{1/x^2}{1/x^2}\)

OpenStudy (anonymous):

oh yeah, the is like the hoz asympt?

OpenStudy (amistre64):

whatever simplifies and still has an 'x' under it; get very very very tiny, for example: 1 1 --- ;-- ------------ = .00000...000001 , a very tiny number right? x 100000...00000

OpenStudy (amistre64):

it is the horizontal asymptote yes

OpenStudy (anonymous):

ok, thx